X=INT(SQR(2)). Which of the following is the value of X in the BASIC program?

a. 1.414
b. 4
c. 0
d. 1
1
Expert's answer
2015-03-03T02:13:38-0500
Answer: b. 4
X = INT( SQR (Y))
This expression meansthat X is equal to the integer part of the square number Y.
Thus, X= INT(SQR(2))=INT(22)=4